OpenStudy (anonymous):

i need to find 3 answers

OpenStudy (anonymous):

and we'll probably need the 3 problems. Wouldn't you agree? :)

OpenStudy (anonymous):

i need to find the equivalent fraction for 2/10 and i need 3 answer for it

OpenStudy (anonymous):

2/ 10 = 1/5 = 4/20 = 200/1,000,...